the inverse function of y=e^3x is y=1/3 lnx
1
Expert's answer
2020-04-20T11:12:15-0400

Consider the function "y=e^{3x}"


Replace "x" and "y" to obtain,



"x=e^{3y}"


Now, take log of base "e" both sides to isolate the exponent as,



"ln(x)=ln(e^{3y})"


Use the property of logarithm "ln(e^{a})=a" to extract the right side of equation from log as,



"ln(x)=3y"


Divide both sides by "3" to obtain,



"y=\frac{1}{3}ln(x)"

Therefore, the inverse of the function "y=e^{3x}" is "y=\frac{1}{3}ln(x)". Hence proved.



Also, note that the function and it's inverse is symmetric about the line "y=x" .